# FinWorld: An All-in-One Open-Source Platform for End-to-End Financial AI Research and Deployment

### v1

@@ A.2. LLM Reasoning Dataset
Our LLM reasoning dataset collection process is designed to provide comprehensive coverage of financial reasoning capabilities, drawing inspiration from established datasets such as FinQA (Chen et al., 2021), FinEval (Zhang et al., 2023), CFLUE (Zhu et al., 2024), Salesforce FinEval (Ke et al., 2025), and ConvFinQA (Chen et al., 2022). We systematically collect diverse types of financial reasoning data through multiple channels and sources, ensuring broad coverage of financial knowledge domains and reasoning patterns.
@@ Appendix B FinWorld as a Visualization Toolkit
Effective visualization is crucial for understanding complex financial AI systems, encompassing datasets, algorithms, models, performance metrics, experimental results and so on. To support comprehensive research and analysis across the entire financial AI workflow, we have developed a comprehensive suite of visualization tools that provide intuitive and interactive representations of our platform’s capabilities. These tools enable researchers to quickly grasp the scope and characteristics of our datasets, understand algorithm performance and model behaviors, analyze training dynamics and convergence patterns, compare different approaches across multiple metrics, and visualize complex financial decision-making processes. The visualization framework facilitates informed decision-making in model selection, hyperparameter tuning, performance evaluation, and research design across all four core financial AI tasks. The visualization tools are implemented using Python, LaTex, Pyecharts (Team, 2025c), and other visualization libraries.
@@ B.2. Performance Visualization
Trading Curve Tools. The proposed module replicates the trading-curve display of FinAgent (Zhang et al., 2024b) while focusing on a single, essential feature. It renders an interactive curve that tracks cumulative return and overlays each model decision with a clearly labelled BUY or SELL marker. An accompanying block of interpretable text records the reasoning behind every trade, summarising the key signals and risk constraints that guided the action. This streamlined presentation allows researchers to relate model logic directly to observed performance, offering a transparent basis for auditing algorithmic behaviour and evaluating how trading decisions drive the evolution of returns over time.
@@ C.1. Factor Model
FinWorld provides comprehensive support for factor model, encompassing both traditional statistical approaches and advanced deep learning methodologies. For traditional statistical factors, we primarily utilize the Alpha158. For deep learning dynamic factor models, we implement state-of-the-art approaches including FactorVAE (Duan et al., 2022) and HireVAE (Wei et al., 2023). FactorVAE leverages variational autoencoders to learn latent factor representations that capture complex market dynamics and temporal dependencies. HireVAE extends this framework with hierarchical structures to model multi-scale temporal patterns in financial data. Notably, the factor model STORM (Zhao et al., 2024) is developed based on this framework, demonstrating the platform’s capability to support cutting-edge research and development in financial AI.
@@ C.2. Financial Market Representation Learning
FinWorld implements comprehensive financial market representation learning capabilities, focusing on stock pool-based representation learning approaches. The platform supports various state-of-the-art representation learning methods, including VAE (Variational Autoencoder) (Kingma et al., 2013), MAE (Masked Autoencoder) (He et al., 2022), and VQVAE (Vector Quantized Variational Autoencoder) (Van Den Oord et al., 2017) for capturing complex market dynamics and relationships. These representation learning methods enable the development of sophisticated market embeddings that capture both local and global patterns in financial data: VAE captures smooth state transitions via continuous latents, MAE learns robust features through masked reconstruction, and VQVAE models discrete regimes with categorical representations.
@@ D.1. Time Series Forecasting
Methods. We evaluate several state-of-the-art ML-based and DL-based time series forecasting models, including: i) ML-based methods: LightGBM (Yang et al., 2020) and XGBoost (Yang et al., 2020); ii) DL-based methods: Autoformer (Wu et al., 2021), Crossformer (Zhang and Yan, 2023), ETSformer (Woo et al., 2022), DLinear (Zeng et al., 2023b), TimesNet (Wu et al., 2022), PatchTST (Nie et al., 2022), TimeMixer (Wang et al., 2024b), and TimeXer (Wang et al., 2024a). These models are selected based on their effectiveness in capturing complex financial time series patterns and their ability to handle high-dimensional data.
Implementation Details. All experiments are conducted on 2 NVIDIA H100 GPUs. Following the implementation practices of Qlib (Yang et al., 2020) and TSLib (Team, 2025e), we adopt consistent architectural configurations across different models while preserving their unique characteristics. For dataset, we utilize a multi-GPU parallel dataloader with a batch size of 64. Each batch contains sequences of 64 historical days and 32 future days, meaning the model observes the past 64 days and predicts the returns over the next 32 days. For network architecture, we use a standardized setup with encoder depth of 4 and decoder depth of 4, where the decoder typically consists of a single MLP layer for output mapping unless the model has specific decoder requirements. The embedding dimension is set to 128 across all models. For optimization, we employ AdamW optimizer with a learning rate of 1e-5 and implement a warmup scheduler for learning rate adjustment, with maximum training epochs of 2,000 and warmup epochs of 200. The loss function is Mean Squared Error (MSE). Additional model-specific configurations are applied based on each model’s architectural requirements and design principles. The reported results are averaged over three runs with different random seeds.
@@ D.2. Algorithmic Trading
DD $\downarrow$ $\mathrm{DD}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\min(\mathrm{r}_{t}-r_{f},0)^{2}}\times\sqrt{N}$ Downside Deviation; annualized standard deviation of returns below the risk-free rate, focusing on downside risk.
Note: $\mathrm{rets}=[r_{1},r_{2},\ldots,r_{T}]$ is the return series, where $r_{t}$ is the return on day $t$; $N$ is the number of periods per year; $r_{f}$ is the risk-free rate.
Methods. We systematically evaluate several Rule-based, ML-based, DL-based, and RL-based methods, including i) Rule-based methods: BUY&HOLD, MACD; ii) ML-based Methods: LightGBM, XGBoost; iii) DL-based Methods: Transformer (Vaswani et al., 2017), LSTM, DLinear (Zeng et al., 2023b); iv) RL-based Methods: PPO (Schulman et al., 2017), SAC (Haarnoja et al., 2018). These models represent a diverse set of approaches commonly used in quantitative finance research and practice.
@@ D.3. Portfolio Management
[…]
@@ E.1. Background
LLMs for Financial AI. The integration of LLMs into financial decision-making processes has fundamentally transformed the landscape of financial AI, giving rise to two predominant developmental trajectories. The first, traditional paradigm leverages large-scale pre-training on generic corpora, followed by supervised fine-tuning (SFT) on domain-specific datasets, a process that has enabled the creation of financial-domain LLMs such as FinBERT (Araci, 2019), FLANG (Shah et al., 2022), and FinGPT (Liu et al., 2023). These models excel at classic financial NLP tasks including sentiment analysis, news classification, and financial document understanding, benefiting from the inclusion of specialized financial terminology and context during pre-training. For example, BloombergGPT (Wu et al., 2023) introduced domain-adapted tokenization and large-scale financial corpus curation, enabling robust out-of-domain generalization while maintaining strong in-domain accuracy on a wide range of financial tasks. Similarly, FinQA (Chen et al., 2021) targets numerical reasoning within financial texts, highlighting the importance of reasoning abilities in addition to conventional text understanding.
@@
The second, more recent development path centers on enhancing the reasoning capabilities of LLMs using RL, inspired by breakthroughs such as DeepSeek-R1 (Guo et al., 2025), which demonstrated substantial improvements in complex reasoning and decision-making tasks. In the context of financial AI, RL-based approaches have facilitated the creation of models capable of more sophisticated financial reasoning, including complex question answering (QA) and multi-step analysis over noisy and unstructured market data. State-of-the-art systems such as Fin-R1 (Liu et al., 2025), Fin-o1 (Qian et al., 2025), and Dianjin-R1 (Zhu et al., 2025) typically adopt a two-stage pipeline: they first apply SFT on curated financial QA datasets to capture domain knowledge, then employ RL to further optimize the models’ reasoning and generalization abilities with task-specific reward functions on challenging financial benchmarks. Notably, this line of work has shown that RL fine-tuning can help mitigate some limitations of SFT-only models, such as shallow pattern matching or overfitting to specific data distributions. However, it is also evident that LLM-based methods remain challenged by issues such as the absence of explicit sequential decision-making mechanisms, high computational overhead (particularly with RL-based training), and an inability to robustly handle the non-stationarity and regime shifts that typify real-world financial markets. Moreover, most current benchmarks still emphasize static QA or single-step predictions, rather than end-to-end financial decision-making in dynamic environments.
@@
LLM Agents for Financial AI. To overcome the intrinsic limitations of static LLM-based models in handling sequential and interactive decision-making tasks, recent research has shifted towards constructing LLM-powered financial agents equipped with advanced agentic mechanisms. These mechanisms, such as external memory, dynamic tool use, self-reflection, and collaborative reasoning, enable LLMs to act not just as passive information processors, but as proactive, context-aware agents capable of adaptive and continual learning. Early efforts, including FinMem (Yu et al., 2024a) and FinRobot (Zhou et al., 2024), demonstrated that the addition of structured memory modules and chain-of-thought reasoning could substantially boost single-asset trading performance and stability, particularly in volatile or data-scarce settings. These systems typically maintain layered or episodic memories that allow agents to recall past market events, dynamically update strategies, and profile assets or markets over time.
@@
Building on these foundations, more recent agentic systems such as FinAgent (Zhang et al., 2024b) have introduced multimodal reasoning capabilities, allowing LLMs to jointly process textual news, structured numerical data, and even K-line (candlestick) chart images via tool-augmented and dual-reflection architectures. This multimodal approach is shown to yield substantial profit gains and improved generalization across diverse datasets. Meanwhile, there has been a pronounced trend toward collaborative and hierarchical multi-agent systems, inspired by real-world financial organizations. For instance, FinCon (Yu et al., 2024b) adopts a manager–analyst architecture, combining conceptual verbal reinforcement and episodic self-critique to facilitate robust sequential decision-making in both single-stock and multi-asset (portfolio) trading. Similarly, TradingAgents (Xiao et al., 2024) simulate an entire trading firm with specialized agents (e.g., fundamental, technical, sentiment, and risk analysts), whose structured debate and coordination have been shown to improve both returns and risk control.
@@
Recent advances have further explored the combination of RL with agentic LLMs to enable dynamic exploration and continual adaptation in non-stationary environments (Feng et al., 2025; Ouyang et al., 2025). For example, group-based RL finetuning enables multiple LLM agents to interact, learn, and adapt strategies in simulated market environments, fostering resilience against sudden market changes. Beyond performance, there is an increasing focus on interpretability and safety in financial LLM agents, with research exploring techniques such as role assignment, self-consistency checking, and real-time human-in-the-loop supervision. Despite these advances, key open challenges remain: efficiently scaling agentic LLMs, mitigating hallucinations and overfitting, improving sample efficiency in RL settings, and reliably deploying such systems in the face of adversarial and ever-evolving market conditions.
@@ E.2. Reinforcement Learning for LLMs
For our financial reasoning tasks, we employ Group Relative Policy Optimization (GRPO) (Shao et al., 2024) as the primary training methodology. GRPO is specifically designed for training large language models in reasoning-intensive scenarios, making it particularly well-suited for financial question answering and document analysis tasks. The algorithm enhances the model’s capability to generate accurate and well-reasoned financial responses through group-based reward normalization and policy optimization.
@@
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\mathrm{old}}}(o_{i,t}\mid q,o_{i,<t})}$ is the importance sampling ratio, and $\epsilon$ is the clipping parameter. Following the approach in DAPO (Yu et al., 2025), we omit the KL divergence regularization term as it has been shown to be unnecessary for chain-of-thought reasoning scenarios.
Following the design of DeepSeek-R1 (Guo et al., 2025), our reward function consists of two main components: format reward and accuracy reward, designed to encourage both proper reasoning structure and correct answers.
@@
Methods. We evaluate our model FinReasoner against open-source LLMs on the test dataset. We systematically evaluate several open-source LLM methods, including DeepSeek-R1 (Guo et al., 2025), Qwen3-8B (Yang et al., 2025), Fin-R1-7B (Guo et al., 2025), and Qwen2.5-7B-Instruct (Team, 2025d). We use the same evaluation protocol as the original paper to ensure fair comparison.
Implementation Details. Our model, FinReasoner, is built on the Qwen-8B language model as the base LLM. In our implementation, we integrate Verl 0.5.0 (Team, 2025f) as an external dependency library within the libs directory. We customize both the trainer and the reward computation as plugins, which are injected into Verl’s trainer (details can be found in the mverl module of our codebase). This implementation approach offers significant advantages: since Verl is frequently updated, the plugin injection method provides greater flexibility for maintaining synchronization with the official Verl library.
@@
Following the design principles of RAGEN (Wang et al., 2025), our reward function in the single-stock trading environment consists of two main components: format reward and trading reward, which are designed to ensure standardized action outputs and effective trading decisions.
@@
Methods. We evaluate our trading models using a simplified FinAgent-based AI Agent framework developed in this work. This framework supports both commercial models (e.g., GPT-4.1, Claude-4-Sonnet) and local models (e.g., Qwen), allowing the underlying LLM to be flexibly replaced. Therefore, our comparison includes DeepSeek-R1 (Guo et al., 2025), Qwen3-8B (Yang et al., 2025), Fin-R1-7B (Guo et al., 2025), Qwen2.5-7B-Instruct (Team, 2025d), GPT-4.1, Claude-4-Sonnet, as well as our own FinReasoner, all serving as the backbone LLMs for the FinAgent. It is worth noting that a detailed description of our FinAgent (LLMs Agents) implementation will be provided in the next section; please refer to Appendix E.3 for details.

### v2

@@ A.2. LLM Reasoning Dataset
Our LLM reasoning dataset collection process is designed to provide comprehensive coverage of financial reasoning capabilities, drawing inspiration from established datasets such as FinQA (chen2021finqa), FinEval (zhang2023fineval), CFLUE (zhu2024cflue), Salesforce FinEval (ke2025demystifyingdomainadaptiveposttrainingfinancial), and ConvFinQA (chen2022convfinqa). We systematically collect diverse types of financial reasoning data through multiple channels and sources, ensuring broad coverage of financial knowledge domains and reasoning patterns.
@@ Appendix B FinWorld as a Visualization Toolkit
Effective visualization is crucial for understanding complex financial AI systems, encompassing datasets, algorithms, models, performance metrics, experimental results and so on. To support comprehensive research and analysis across the entire financial AI workflow, we have developed a comprehensive suite of visualization tools that provide intuitive and interactive representations of our platform’s capabilities. These tools enable researchers to quickly grasp the scope and characteristics of our datasets, understand algorithm performance and model behaviors, analyze training dynamics and convergence patterns, compare different approaches across multiple metrics, and visualize complex financial decision-making processes. The visualization framework facilitates informed decision-making in model selection, hyperparameter tuning, performance evaluation, and research design across all four core financial AI tasks. The visualization tools are implemented using Python, LaTex, Pyecharts (pyecharts), and other visualization libraries.
@@ B.2. Performance Visualization
Trading Curve Tools. The proposed module replicates the trading-curve display of FinAgent (zhang2024multimodal) while focusing on a single, essential feature. It renders an interactive curve that tracks cumulative return and overlays each model decision with a clearly labelled BUY or SELL marker. An accompanying block of interpretable text records the reasoning behind every trade, summarising the key signals and risk constraints that guided the action. This streamlined presentation allows researchers to relate model logic directly to observed performance, offering a transparent basis for auditing algorithmic behaviour and evaluating how trading decisions drive the evolution of returns over time.
@@ C.1. Factor Model
FinWorld provides comprehensive support for factor model, encompassing both traditional statistical approaches and advanced deep learning methodologies. For traditional statistical factors, we primarily utilize the Alpha158. For deep learning dynamic factor models, we implement state-of-the-art approaches including FactorVAE (duan2022factorvae) and HireVAE (wei2023hirevae). FactorVAE leverages variational autoencoders to learn latent factor representations that capture complex market dynamics and temporal dependencies. HireVAE extends this framework with hierarchical structures to model multi-scale temporal patterns in financial data. Notably, the factor model STORM (zhao2024storm) is developed based on this framework, demonstrating the platform’s capability to support cutting-edge research and development in financial AI.
@@ C.2. Financial Market Representation Learning
FinWorld implements comprehensive financial market representation learning capabilities, focusing on stock pool-based representation learning approaches. The platform supports various state-of-the-art representation learning methods, including VAE (Variational Autoencoder) (vae), MAE (Masked Autoencoder) (mae), and VQVAE (Vector Quantized Variational Autoencoder) (vqvae) for capturing complex market dynamics and relationships. These representation learning methods enable the development of sophisticated market embeddings that capture both local and global patterns in financial data: VAE captures smooth state transitions via continuous latents, MAE learns robust features through masked reconstruction, and VQVAE models discrete regimes with categorical representations.
@@ D.1. Time Series Forecasting
Methods. We evaluate several state-of-the-art ML-based and DL-based time series forecasting models, including: i) ML-based methods: LightGBM (yang2020qlib) and XGBoost (yang2020qlib); ii) DL-based methods: Autoformer (wu2021autoformer), Crossformer (zhang2023crossformer), ETSformer (woo2022etsformer), DLinear (zeng2023transformers), TimesNet (wu2022timesnet), PatchTST (nie2022time), TimeMixer (wang2024timemixer), and TimeXer (wang2024timexer). These models are selected based on their effectiveness in capturing complex financial time series patterns and their ability to handle high-dimensional data.
Implementation Details. All experiments are conducted on 2 NVIDIA H100 GPUs. Following the implementation practices of Qlib (yang2020qlib) and TSLib (tslib), we adopt consistent architectural configurations across different models while preserving their unique characteristics. For dataset, we utilize a multi-GPU parallel dataloader with a batch size of 64. Each batch contains sequences of 64 historical days and 32 future days, meaning the model observes the past 64 days and predicts the returns over the next 32 days. For network architecture, we use a standardized setup with encoder depth of 4 and decoder depth of 4, where the decoder typically consists of a single MLP layer for output mapping unless the model has specific decoder requirements. The embedding dimension is set to 128 across all models. For optimization, we employ AdamW optimizer with a learning rate of 1e-5 and implement a warmup scheduler for learning rate adjustment, with maximum training epochs of 2,000 and warmup epochs of 200. The loss function is Mean Squared Error (MSE). Additional model-specific configurations are applied based on each model’s architectural requirements and design principles. The reported results are averaged over three runs with different random seeds.
@@ D.2. Algorithmic Trading
$D$ $\downarrow$ $\mathrm{DD}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\min(\mathrm{r}_{t}-r_{f},0)^{2}}\times\sqrt{N}$ Downside Deviation; annualized standard deviation of returns below the risk-free rate, focusing on downside risk.
Note: $\mathrm{rets}=[r_{1},r_{2},\ldots,r_{T}]$ is the return series, where $r_{t}$ is the return on day $t$; NN is the number of periods per year; $r_{f}$ is the risk-free rate.
Methods. We systematically evaluate several Rule-based, ML-based, DL-based, and RL-based methods, including i) Rule-based methods: BUY&HOLD, MACD; ii) ML-based Methods: LightGBM, XGBoost; iii) DL-based Methods: Transformer (vaswani2017attention), LSTM, DLinear (zeng2023transformers); iv) RL-based Methods: PPO (schulman2017proximal), SAC (haarnoja2018soft). These models represent a diverse set of approaches commonly used in quantitative finance research and practice.
@@ D.3. Portfolio Management
Methods. We systematically evaluate several ML-based, DL-based, and RL-based methods. Since we can apply the top-$k$ dropout strategy  (yang2020qlib) after any ML&DL-based forecasting model to construct a portfolio, our ML&DL-based methods include all ML&DL-based forecasting models. Specifically, we consider: i) Rule-based methods: BUY&HOLD; ii) ML-based Methods: LightGBM, XGBoost; iii) DL-based Methods: Autoformer (wu2021autoformer), Crossformer (zhang2023crossformer), ETSformer (woo2022etsformer), DLinear (zeng2023transformers), TimesNet (wu2022timesnet), PatchTST (nie2022time), TimeMixer (wang2024timemixer), and TimeXer (wang2024timexer); iv) RL-based Methods: PPO (schulman2017proximal), SAC (haarnoja2018soft). These models represent a diverse set of approaches commonly used in quantitative finance research and practice.
Implementation Details. All experiments are conducted on 2 NVIDIA H100 GPUs. For all DL-based methods, we use similar dataset splits, network architectures, and optimizer configurations as in the forecasting task. In addition, all ML&DL-based methods use the Top-$k$ Dropout strategy (yang2020qlib) for portfolio construction. For RL-based methods, the initial cash in the environment is set to $1\times 10^{5}$, and the transaction fee rate is $1\times 10^{-4}$. We use 4 parallel environments for data sampling, with a sampling step size of 512. The policy mini-batch size is set to 128 and the value mini-batch size to 16. The optimizer used is AdamW, with the policy network learning rate set to $1\times 10^{-5}$ and the value network learning rate set to $1\times 10^{-6}$. The maximum number of training steps is $1\times 10^{8}$, and the learning rate warmup is applied starting from step 0. Additional model-specific configurations are applied based on each model’s architectural requirements and design principles.The reported results are averaged over three runs with different random seeds.
@@ E.1. Background
LLMs for Financial AI. The integration of LLMs into financial decision-making processes has fundamentally transformed the landscape of financial AI, giving rise to two predominant developmental trajectories. The first, traditional paradigm leverages large-scale pre-training on generic corpora, followed by supervised fine-tuning (SFT) on domain-specific datasets, a process that has enabled the creation of financial-domain LLMs such as FinBERT (araci2019finbert), FLANG (shah2022flang), and FinGPT (liu2023fingpt). These models excel at classic financial NLP tasks including sentiment analysis, news classification, and financial document understanding, benefiting from the inclusion of specialized financial terminology and context during pre-training. For example, BloombergGPT (wu2023bloomberggpt) introduced domain-adapted tokenization and large-scale financial corpus curation, enabling robust out-of-domain generalization while maintaining strong in-domain accuracy on a wide range of financial tasks. Similarly, FinQA (chen2021finqa) targets numerical reasoning within financial texts, highlighting the importance of reasoning abilities in addition to conventional text understanding.
@@
The second, more recent development path centers on enhancing the reasoning capabilities of LLMs using RL, inspired by breakthroughs such as DeepSeek-R1 (guo2025deepseek), which demonstrated substantial improvements in complex reasoning and decision-making tasks. In the context of financial AI, RL-based approaches have facilitated the creation of models capable of more sophisticated financial reasoning, including complex question answering (QA) and multi-step analysis over noisy and unstructured market data. State-of-the-art systems such as Fin-R1 (liu2025fin), Fin-o1 (qian2025fino1), and Dianjin-R1 (zhu2025dianjin) typically adopt a two-stage pipeline: they first apply SFT on curated financial QA datasets to capture domain knowledge, then employ RL to further optimize the models’ reasoning and generalization abilities with task-specific reward functions on challenging financial benchmarks. Notably, this line of work has shown that RL fine-tuning can help mitigate some limitations of SFT-only models, such as shallow pattern matching or overfitting to specific data distributions. However, it is also evident that LLM-based methods remain challenged by issues such as the absence of explicit sequential decision-making mechanisms, high computational overhead (particularly with RL-based training), and an inability to robustly handle the non-stationarity and regime shifts that typify real-world financial markets. Moreover, most current benchmarks still emphasize static QA or single-step predictions, rather than end-to-end financial decision-making in dynamic environments.
@@
LLM Agents for Financial AI. To overcome the intrinsic limitations of static LLM-based models in handling sequential and interactive decision-making tasks, recent research has shifted towards constructing LLM-powered financial agents equipped with advanced agentic mechanisms. These mechanisms, such as external memory, dynamic tool use, self-reflection, and collaborative reasoning, enable LLMs to act not just as passive information processors, but as proactive, context-aware agents capable of adaptive and continual learning. Early efforts, including FinMem (yu2024finmem) and FinRobot (zhou2024finrobot), demonstrated that the addition of structured memory modules and chain-of-thought reasoning could substantially boost single-asset trading performance and stability, particularly in volatile or data-scarce settings. These systems typically maintain layered or episodic memories that allow agents to recall past market events, dynamically update strategies, and profile assets or markets over time.
@@
Building on these foundations, more recent agentic systems such as FinAgent (zhang2024multimodal) have introduced multimodal reasoning capabilities, allowing LLMs to jointly process textual news, structured numerical data, and even K-line (candlestick) chart images via tool-augmented and dual-reflection architectures. This multimodal approach is shown to yield substantial profit gains and improved generalization across diverse datasets. Meanwhile, there has been a pronounced trend toward collaborative and hierarchical multi-agent systems, inspired by real-world financial organizations. For instance, FinCon (yu2024fincon) adopts a manager–analyst architecture, combining conceptual verbal reinforcement and episodic self-critique to facilitate robust sequential decision-making in both single-stock and multi-asset (portfolio) trading. Similarly, TradingAgents (xiao2024tradingagents) simulate an entire trading firm with specialized agents (e.g., fundamental, technical, sentiment, and risk analysts), whose structured debate and coordination have been shown to improve both returns and risk control.
@@
Recent advances have further explored the combination of RL with agentic LLMs to enable dynamic exploration and continual adaptation in non-stationary environments (feng2025group; agent-r1). For example, group-based RL finetuning enables multiple LLM agents to interact, learn, and adapt strategies in simulated market environments, fostering resilience against sudden market changes. Beyond performance, there is an increasing focus on interpretability and safety in financial LLM agents, with research exploring techniques such as role assignment, self-consistency checking, and real-time human-in-the-loop supervision. Despite these advances, key open challenges remain: efficiently scaling agentic LLMs, mitigating hallucinations and overfitting, improving sample efficiency in RL settings, and reliably deploying such systems in the face of adversarial and ever-evolving market conditions.
@@ E.2. Reinforcement Learning for LLMs
For our financial reasoning tasks, we employ Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) as the primary training methodology. GRPO is specifically designed for training large language models in reasoning-intensive scenarios, making it particularly well-suited for financial question answering and document analysis tasks. The algorithm enhances the model’s capability to generate accurate and well-reasoned financial responses through group-based reward normalization and policy optimization.
@@
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\mathrm{old}}}(o_{i,t}\mid q,o_{i,<t})}$ is the importance sampling ratio, and $\epsilon$ is the clipping parameter. Following the approach in DAPO (yu2025dapo), we omit the KL divergence regularization term as it has been shown to be unnecessary for chain-of-thought reasoning scenarios.
Following the design of DeepSeek-R1 (guo2025deepseek), our reward function consists of two main components: format reward and accuracy reward, designed to encourage both proper reasoning structure and correct answers.
@@
Methods. We evaluate our model FinReasoner against open-source LLMs on the test dataset. We systematically evaluate several open-source LLM methods, including DeepSeek-R1 (guo2025deepseek), Qwen3-8B (yang2025qwen3), Fin-R1-7B (guo2025deepseek), and Qwen2.5-7B-Instruct (qwen2025qwen25technicalreport). We use the same evaluation protocol as the original paper to ensure fair comparison.
Implementation Details. Our model, FinReasoner, is built on the Qwen-8B language model as the base LLM. In our implementation, we integrate Verl 0.5.0 (verl) as an external dependency library within the libs directory. We customize both the trainer and the reward computation as plugins, which are injected into Verl’s trainer (details can be found in the mverl module of our codebase). This implementation approach offers significant advantages: since Verl is frequently updated, the plugin injection method provides greater flexibility for maintaining synchronization with the official Verl library.
@@
Following the design principles of RAGEN (ragen), our reward function in the single-stock trading environment consists of two main components: format reward and trading reward, which are designed to ensure standardized action outputs and effective trading decisions.
@@
Methods. We evaluate our trading models using a simplified FinAgent-based AI Agent framework developed in this work. This framework supports both commercial models (e.g., GPT-4.1, Claude-4-Sonnet) and local models (e.g., Qwen), allowing the underlying LLM to be flexibly replaced. Therefore, our comparison includes DeepSeek-R1 (guo2025deepseek), Qwen3-8B (yang2025qwen3), Fin-R1-7B (guo2025deepseek), Qwen2.5-7B-Instruct (qwen2025qwen25technicalreport), GPT-4.1, Claude-4-Sonnet, as well as our own FinReasoner, all serving as the backbone LLMs for the FinAgent. It is worth noting that a detailed description of our FinAgent (LLMs Agents) implementation will be provided in the next section; please refer to Appendix E.3 for details.